OpenStudy (anonymous):

Assume that u,v E ( -pi/2, pi/2 ). use mean value theorem to prove that |tanu-tanv| >= |u-v|

OpenStudy (anonymous):

let f(x)=tan x f(x) and f'(x) are continuous on E withot loss of generality let u>v so there are some values v<w<u such that \[\frac{\tan u - \tan v}{u-v}=f'(w)=1+\tan^2 w \ge 1\]

OpenStudy (anonymous):

@mukushla : Could you explain me a little bit more, please.

OpenStudy (anonymous):

@JamesJ @TuringTest help plz my english is not well :(

OpenStudy (anonymous):

@waterineyes

OpenStudy (anonymous):

@FoolAroundMath

OpenStudy (anonymous):

@JamesJ @Hero could you help me with this

OpenStudy (anonymous):

Let f(x) be defined in [a, b ] and differentiable in (a, b). there exists a number c betwween a abd b such that f'(c) = [f(b) - f(a)]/(b-a). That is mean value theorem, Given f(x) = tan (x) the derivative is f'(x) = sec^2 (x) Assume u > v. There exist w such v < w < u and [tan(u) - tan(v)] / (u - v) = sec^2 (x). since sec (x) either >= 1 or =< -1, sec^2 (x) >= 1 (We can also let sec^2 (x) = 1 + tan^2 (x) >= 1) So [tan (u) - tan(v)] >= (u - v) If you don't assume u > v, then we need to use absolute value on both sides.

OpenStudy (anonymous):

THank @Mr._To and @mukushla igot it now